# Multimodal Fusion on Low-quality Data: A Comprehensive Survey

### v1

@@ 5.3 Uncertainty-aware dynamic fusion
Inspired by TMC, researchers deployed SL-based uncertainty-aware multimodal fusion in many scenarios, including medical image classification [149], object detection [148, 157], semantic segmentation [2]. In whole-slide medical image classification task, by integrating evidences from modalities of multiple scales, they significantly improve the medical classification performance in autoimmune diseases detection and Fibroma classification tasks. Similarly, Chen et al. [158] also leverage Dirichlet distribution to model the uncertainty of multimodal representations and then fuse the segmentation and classification representations adaptively.
@@
Although TMC exhibits a great performance, Liu et al. [150] identified its limitations, including inability to guarantee an overall decrease in uncertainty during the fusion of diverse views and neglecting of consistency across views. To address these issues, they proposed an improved opinion aggregation framework, ensuring consistency across views and achieving more reliable predictions. Xu et al. [159] pointed out that an important characteristic of TMC is ”After integrating another opinion into the original opinion, the obtained uncertainty mass will be reduced”. They argued this since when incorporating an unreliable or conflicting opinion, the uncertainty should increase. To solve this, they proposed a conflictive opinion aggregation strategy and theoretically prove the uncertainty would increase for conflictive instances.
@@
Entropy is the most natural and straightforward way to capture modal-wise uncertainty. Tian et al. [144] proposed an entropy-based multimodal fusion method, named Uncertainty-aware Noisy-Or Multimodal Fusion (UNO). Specifically, UNO first introduces several uncertainty measurements, including predictive entropy, mutual information based on Monte Carlo Dropout, as well as deterministic entropy. It then proceeds to combine these uncertainty metrics for fusion by selecting the most conservative (highest uncertainty) measurement and weighting the outputs from each modalities based on the uncertainty. Finally, a simple but effective noisy-or fusion scheme is deployed to fusing decisions from multiple modalities. In multimodal semantic segmentation task, UNO can be used to improve the robustness in the presence of various unknown input degradation.
@@
Gaussian distributions are another commonly used way to model the uncertainty. For example, a multivariate normal distribution can be used to model the distribution of multimodal feature. The variance of estimated distribution are capable of representing modality-wise uncertainty. Base on this point, Geng et al. [151] proposed a novel unsupervised multimodal learning method, called Dynamic Uncertainty-Aware Network (DUA-Net). Integrating multiple modalities of noisy data in unsupervised settings is a challenging task. Traditional multimodal methods either treat each modality as equally important, or adjust the weights of different modalities to fixed values, which fail to capture the dynamic uncertainty in multimodal data. By integrating the intrinsic information from multiple modalities under the guidance of uncertainty estimates from a generative perspective, DUA-Net obtains noise-free representations. With the help of uncertainty estimates, DUA-Net can weight each modality of a single sample according to its data quality, thus fully utilizing high-quality samples or modalities while mitigating the influence of noisy samples or modalities. The proposed method is shown to be effective in various unsupervised multimodal learning tasks. For audio-visual emotion recognition, Tellamekala et al. [152] proposed Calibrated and Ordinal Latent Distributions (COLD) to model uncertainty for each modality. In this work, the variance of high-dimensional latent distributions is learned first for each modality as a measure of uncertainty. In object detection task, previous work [146] proposed a probabilistic ensemble technique (ProbEnsemble) to combine the predictions from multiple detectors, each using different input modalities and architectures. In this work, the authors took the form of a Gaussian with a single variance to measure the uncertainty of unimodal predictions. Then, they probabilistically fused outputs from multiple modalities follow Bayesian’ rule. The proposed probabilistic ensemble framework can estimate the likelihood of each detection proposal and then combine them into a more accurate and robust detection result.
@@
Normal-inverse Gamma distributions hierarchically characterizes the uncertainties with parameters of high-order distributions. In the field of multimodal regression, a recent paper [153] presents a new multimodal fusion approach based on a Normal-Inverse Gaussian (NIG) mixture model. This NIG mixture model is similar to the decision-level fusion method, which combines multiple NIG distributions to output the uncertainty of each modality and the overall uncertainty. In this work, the NIG distribution is deployed to model modal-specific epistemic uncertainty (EU) and aleatoric uncertainty (AU). The contribution of each modality to the final prediction can be dynamically adjusted based on its quality and uncertainty, resulting in more reliable and accurate prediction. The proposed method is implemented using expert decision fusion and achieves significant improvements in both accuracy and uncertainty estimation in multimodal regression tasks.
@@
Prediction confidence methods are designed to directly output the confidence. Different from the uncertainty estimation algorithms, confidence calibration methods aim to obtain confidence by directly calibrating the classification results without modeling the unknown data distributions. For example, the Maximum Class Probability (MCP) [160] can be regarded as a baseline method to obtain prediction confidence. Although effective in classification, MCP usually leads to over-confidence especially for erroneous prediction. Therefore, the True-Class Probability (TCP) [161] is employed to obtain more reliable classification confidence. Han et al. [154] proposed a multimodal classification method called Multimodal Dynamics that evaluates both feature-level and modality-level informativeness to selectively fuse multiple modalities in a trustworthy manner. Specifically, it uses a sparse gating strategy and true class probability estimation to assess the varying importance of features and modalities for each sample, achieving more reliable fusion when data quality varies across samples and modalities. Specifically, a modality-level gating strategy is employed to perform dynamic fusion as
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation operator, $\mathbf{h}^{1},\cdots,\mathbf{h}^{M}$ and $\mathbf{h}$ are unimodal and multimodal representations respectively. $conf^{m}$ is the predicted confidence for the $m$-th modality which measures the informativeness. Zhang et al. [33] proposed a confidence-aware fusion method to selectively integrate features from RGB and thermal modalities for multi-spectral pedestrian detection. It predicts the classification confidence of each modality, then performs feature re-weighting to emphasize cues from more reliable modality as well as suppress the less useful one based on the confidence scores.
Theoretical justification has been proposed to shed light upon the advantage of uncertainty-aware dynamic multimodal learning. In multimodal learning theory literature, seminal work shows multimodal learning is provably better than unimodal from the perspective of generalization ability [162]. Similarly, recent work [145] illustrates that uncertainty-aware dynamic multimodal learning can establish a tighter generalization error bound comparing to traditional static methods. Considering ensemble-like late fusion strategy in classification setting, we can perform uncertainty-aware weighting fusion in decision-level according to the following rule
@@ 5.4 Discussion
Dynamic multimodal learning methods focuses on the variance of modality quality with respect to sample, time or space, which widely exists but tends to be neglected. The dynamic fusion methods include heuristic (mostly designed for specific applications), attention-based (typically for representation fusion) and uncertainty-aware (modeling the modality and sample level uncertainty for fusion) strategies. There are great potentials for dynamic multimodal learning. First, the dynamic principle could be considered in SOTA multimodal models, e.g., CLIP. Second, there are lots of dynamic scenarios in real applications (e.g., autonomous driving, medical image fusion), so designing application-specific dynamic fusion strategies is interesting. For example, in multimodal medical images, we can dynamically fuse them at path-level, which may providing better flexibility and interpretablility.
@@ 6 Conclusion
In this survey, we comprehensively investigate the main challenges of multimodal learning in the wild: noisy, partial, unbalanced, and quality-varying multimodal data. Some of them such as noisy multimodal data have been studied for a long time, but more recent interest in representation and translation have led to a large number of new multimodal algorithms and exciting multimodal applications. For example, low-quality multimodal data for foundation models, e.g., CLIP. (2) low-quality multimodal data for other tasks beyond fusion, e.g., alignment (retrieval). (3) Other low-quality forms beyond noise, incomplete, unbalanced, and dynamic (e.g., adversarial). We believe that our survey will help catalog future research papers and better understand the remaining unresolved problems facing multimodal machine learning.

### v2

@@ 5.3 Uncertainty-aware dynamic fusion
Inspired by TMC, researchers deployed SL-based uncertainty-aware multimodal fusion in many scenarios, including medical image classification [149], object detection [148, 157], semantic segmentation [2]. In whole-slide medical image classification task, by integrating evidences from modalities of multiple scales, they significantly improve the medical classification performance in autoimmune diseases detection and Fibroma classification tasks. Similarly, Chen et al[158] also leverage Dirichlet distribution to model the uncertainty of multimodal representations and then fuse the segmentation and classification representations adaptively.
@@
Although TMC exhibits a great performance, Liu et al[150] identified its limitations, including inability to guarantee an overall decrease in uncertainty during the fusion of diverse views and neglecting of consistency across views. To address these issues, they proposed an improved opinion aggregation framework, ensuring consistency across views and achieving more reliable predictions. Xu et al[159] pointed out that an important characteristic of TMC is ”After integrating another opinion into the original opinion, the obtained uncertainty mass will be reduced”. They argued this since when incorporating an unreliable or conflicting opinion, the uncertainty should increase. To solve this, they proposed a conflictive opinion aggregation strategy and theoretically proved the uncertainty would increase for conflictive instances.
@@
Entropy is the most natural and straightforward way to capture modal-wise uncertainty. Tian et al[144] proposed an entropy-based multimodal fusion method, named Uncertainty-aware Noisy-Or Multimodal Fusion (UNO). Specifically, UNO first introduces several uncertainty measurements, including predictive entropy, mutual information based on Monte Carlo Dropout, as well as deterministic entropy. It then proceeds to combine these uncertainty metrics for fusion by selecting the most conservative (highest uncertainty) measurement and weighting the outputs from each modalities based on the uncertainty. Finally, a simple but effective noisy-or fusion scheme is deployed to fusing decisions from multiple modalities. In multimodal semantic segmentation task, UNO can be used to improve the robustness in the presence of various unknown input degradation.
@@
Gaussian distributions are another commonly used way to model the uncertainty. For example, a multivariate normal distribution can be used to model the distribution of multimodal feature. The variance of estimated distribution are capable of representing modality-wise uncertainty. Based on this point, Geng et al[151] proposed a novel unsupervised multimodal learning method, called Dynamic Uncertainty-Aware Network (DUA-Net). Integrating multiple modalities of noisy data in unsupervised settings is a challenging task. Traditional multimodal methods either treat each modality as equally important, or adjust the weights of different modalities to fixed values, which fail to capture the dynamic uncertainty in multimodal data. By integrating the intrinsic information from multiple modalities under the guidance of uncertainty estimates from a generative perspective, DUA-Net obtains noise-free representations. With the help of uncertainty estimates, DUA-Net can weight each modality of a single sample according to its data quality, thus fully utilizing high-quality samples or modalities while mitigating the influence of noisy samples or modalities. The proposed method is shown to be effective in various unsupervised multimodal learning tasks. For audio-visual emotion recognition, Tellamekala et al[152] proposed Calibrated and Ordinal Latent Distributions (COLD) to model uncertainty for each modality. In this work, the variance of high-dimensional latent distributions is learned first for each modality as a measure of uncertainty. In object detection task, previous work [146] proposed a probabilistic ensemble technique (ProbEnsemble) to combine the predictions from multiple detectors, each using different input modalities and architectures. In this work, the authors took the form of a Gaussian with a single variance to measure the uncertainty of unimodal predictions. Then, they probabilistically fused outputs from multiple modalities following Bayes’ rule. The proposed probabilistic ensemble framework can estimate the likelihood of each detection proposal and then combine them into a more accurate and robust detection result.
@@
Normal-inverse Gamma distribution hierarchically characterizes the uncertainties with parameters of high-order distributions. In the field of multimodal regression, a recent paper [153] presents a new multimodal fusion approach based on a Normal-Inverse Gaussian (NIG) mixture model. This NIG mixture model is similar to the decision-level fusion method, which combines multiple NIG distributions to output the uncertainty of each modality and the overall uncertainty. In this work, the NIG distribution is deployed to model modal-specific epistemic uncertainty (EU) and aleatoric uncertainty (AU). The contribution of each modality to the final prediction can be dynamically adjusted based on its quality and uncertainty, resulting in more reliable and accurate prediction. The proposed method is implemented using expert decision fusion and achieves significant improvements in both accuracy and uncertainty estimation in multimodal regression tasks.
@@
Prediction confidence methods are designed to directly output the confidence. Different from the uncertainty estimation algorithms, confidence calibration methods aim to obtain confidence by directly calibrating the classification results without modeling the unknown data distributions. For example, the Maximum Class Probability (MCP) [160] can be regarded as a baseline method to obtain prediction confidence. Although effective in classification, MCP usually leads to over-confidence especially for erroneous prediction. Therefore, the True-Class Probability (TCP) [161] is employed to obtain more reliable classification confidence. Han et al[154] proposed a multimodal classification method called Multimodal Dynamics that evaluates both feature-level and modality-level informativeness to selectively fuse multiple modalities in a trustworthy manner. Specifically, it uses a sparse gating strategy and true class probability estimation to assess the varying importance of features and modalities for each sample, achieving more reliable fusion when data quality varies across samples and modalities. Specifically, a modality-level gating strategy is employed to perform dynamic fusion as
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation operator, $\mathbf{h}^{1},\cdots,\mathbf{h}^{M}$ and $\mathbf{h}$ are unimodal and multimodal representations respectively. $conf^{m}$ is the predicted confidence for the $m$-th modality which measures the informativeness. Zhang et al[33] proposed a confidence-aware fusion method to selectively integrate features from RGB and thermal modalities for multi-spectral pedestrian detection. It predicts the classification confidence of each modality, then performs feature re-weighting to emphasize cues from more reliable modalityies as well as suppress the less useful ones based on the confidence scores.
Theoretical justification has been proposed to shed light upon the advantage of uncertainty-aware dynamic multimodal learning. In multimodal learning theory literature, seminal work shows multimodal learning is provably better than unimodal from the perspective of generalization ability [162]. Similarly, recent work [145] illustrates that uncertainty-aware dynamic multimodal learning can establish a tighter generalization error bound compared to traditional static methods. Combining with ensemble-like late fusion strategy in classification setting, we can perform uncertainty-aware weighting fusion in decision-level according to the following rule
@@ 5.4 Discussion
Dynamic multimodal learning methods focuse on the variance of modality quality with respect to sample, time or space, which widely exists but tends to be neglected. The dynamic fusion methods include heuristic (mostly designed for specific applications), attention-based (typically for representation fusion) and uncertainty-aware (modeling the modality and sample level uncertainty for fusion) strategies. There are great potentials for dynamic multimodal learning. First, the dynamic principle could be considered in SOTA multimodal models, e.g., CLIP. Second, there are lots of dynamic scenarios in real applications (e.g., autonomous driving, medical image fusion), so designing application-specific dynamic fusion strategies is interesting. For example, in multimodal medical images, we can dynamically fuse them at path-level, which may providing better flexibility and interpretablility.
@@ 6 Conclusion
In this survey, we comprehensively investigate the main challenges of multimodal learning in the wild: noisy, partial, unbalanced, and quality-varying multimodal data. Some of them such as noisy multimodal data have been studied for a long time, but more recent interests in representation and translation have led to a large number of new multimodal algorithms and exciting multimodal applications. For example, low-quality multimodal data for foundation models, e.g., CLIP. (2) low-quality multimodal data for other tasks beyond fusion, e.g., alignment (retrieval). (3) Other low-quality forms beyond noise, incomplete, unbalanced, and dynamic (e.g., adversarial). We believe that our survey will help catalog future research papers and better understand the remaining unresolved problems facing multimodal machine learning.